OpenStudy (anonymous):

BEST ANSWER WILL BE REWARDED!!! What is the value of p – (–3.5) when p = 8.2?

OpenStudy (anonymous):

@morganchavez

OpenStudy (anonymous):

8.2 - ( -3.5 ) = 8.2 + 3.5 = 11.7

OpenStudy (anonymous):

Round each number to the nearest tenth. What is the best estimate for the answer to 294.32 – 198.37? A. 90 B. 95.9 C. 96 D. 96.1

OpenStudy (anonymous):

thx do u mind if i ask u more

OpenStudy (anonymous):

Round each number to the nearest tenth. What is the best estimate for the answer to 294.32 – 198.37? A. 90 B. 95.9 C. 96 D. 96.1

OpenStudy (anonymous):

c

OpenStudy (anonymous):

why

OpenStudy (anonymous):

BYE @lol1help

OpenStudy (anonymous):

why is the awnser c

OpenStudy (anonymous):

why u totally r asking questions in my box go make ur own box

OpenStudy (anonymous):

I'll tell you if you're right if you gie the question and your own answer

OpenStudy (anonymous):

@HerbvanH What value of x makes the equation true? –7.12 = –4.8 + x A. –11.92 B. –11.2 C. –2.32 D. 2.32 d?

OpenStudy (anonymous):

-7.12 + 4.8 = x

OpenStudy (anonymous):

is it d